Compare two turntables, A and B. The turntables have the same surface, but turntable A has twice as big radius than turntable B. We can predict that the maximum frequency of rotation at which coin still does not slide off the turntable A is

A) twice as big as the frequency for turntable B

B) same as the frequency for turntable B

C) 1/2 of the frequency of turntable B

D) 21/2 times the frequency of turntable B

E) 2-1/2 time the frequency of turntable B
Answer #1

centripetal acceleration is given by

a = r w^2

for constant acceleration

r is inversely proportional to square of w

so

r1 / r2 = (w2/w1) ^2

r1 / 2 r1 = ( w2/ w1) ^2

w2 = w1/ sqrt (2)

so option (e) is correct

=======

Comment before rate in case any doubt, will reply for sure.. goodluck